X+y2(squared)=4
Determine whether y is a function of x

Answer by Theo(13342)

y is not a function of x.

start with x + y^2 = 4

subtract x from both sides of the equation to get y^2 = 4-x

take the square root of both sides of the equation to get y = plus or minus sqrt(4-x).

you are getting 2 values of y for every value of x.

that means that y is not a function of x.

in order for it to be a function, you must get one and only one value of y for each and every value of x.